40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri May 1 01:49:55 2020, gautam, Update, ALS, ASY commissioning 
    Reply  Fri May 1 22:44:17 2020, gautam, Update, ALS, ASY M2 PZT damaged 
       Reply  Sat May 2 02:35:18 2020, Koji, Update, ALS, ASY M2 PZT damaged 
Message ID: 15316     Entry time: Fri May 1 22:44:17 2020     In reply to: 15315     Reply to this: 15317
Author: gautam 
Type: Update 
Category: ALS 
Subject: ASY M2 PZT damaged 

I went to EY and saw that the HV power supply was only putting out 50 V and had hit the current limit of 10 mA (nominally, it should be 100 V, drawing ~7mA). This is definitely a problem that has come up after the power shutdown event, as when I re-energized the HV power supply at EY, I had confirmed that it was putting out the nominal values (the supply was not labelled with these nominal numbers so I had to label it). Or maybe I broke it while running the dither alignment tests yesterday, even though I never drove the PZTs above 50 Hz with more than 1000cts (= 300 mV * gain 5 in the HV amplifier = 1.5 V ) amplitude.

The problem was confirmed to be with the M2 PZT (YAW channel) and not the electronics by driving the M2 PZT with the M1 channels. Separately, the M1 PZT could be driven by the M2 channels. I also measured the capacitance of the YAW channels and found it to be nearly twice (~7 uF) of the expected 3 uF - this particular PZT is different from the three others in use by the ASX and ASY system, it is an older vintage, so maybe it just failed? 😔 

I don't want to leave 100 V on in this state, so the HV supply at EY was turned off. Good GTRY was recovered by manual alignment of the mirror mounts. If someone has a spare PZT, we can replace it, but for now, we just have to live with manually aligning the green beam often.

Quote:

Could be that the power outage busted something in the drive electronics. 

ELOG V3.1.3-